Consider a population proportion 0.68. The standard error for proportion for n=20 is
1. 0.0109
2. 0.1043
3. 0.2176
4. 0.1404
5. 0.034
1
Expert's answer
2020-11-10T20:00:18-0500

The standard error of a sample proportion can be calculated as:

"\sqrt{\frac{p(1-p)}{n}}"

given: p= 0.68, n= 20


Substitute the values


standard error= "\sqrt{\frac{0.68(1-0.68)}{20}}" = 0.104307